OpenStudy (anonymous):

determine the standard form of the equation of the line that passes throught (-5,0) and (0,-9)

OpenStudy (p0sitr0n):

lets find the slope the slope m is the variation between points on line lets usethe formula \[m =Deltay/Deltax\] so it gives -9/5 y = -9/5x +b to find b, replace x and y by a point 0 = -9/5 *-5 +b b =9 y = -9/5x +9

OpenStudy (p0sitr0n):

\[m = \Delta y/\Delta x\]

OpenStudy (anonymous):

ok thanks..is that standard form?

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. 9x + 5y = -45 --> Standard Form